# Information Shapes Koopman Representation

### v1

@@ Appendix D Limitations and Future Directions
A current limitation of our framework is that it does not address the sample complexity or non-asymptotic convergence of the Koopman representation; future work could explore more rigorous theoretical analyses in this direction. In addition, recent studies have highlighted connections between kernel methods and information theory (Bach, 2022), suggesting an interesting avenue for extending conventional kernel techniques in Koopman theory through an information-theoretic perspective.
@@ Appendix E Key Technical Definitions and Related Properties
Definition E.2 (Effective Dimension (Roy & Vetterli, 2007))
@@ Appendix F Theoretical Framework
Figure 7: The autoregressive structure of Koopman representation. Top row (solid arrows): the original states $x_{t-n}$ evolve under the nonlinear map $T$. Bottom row (dashed arrows): the states are first encoded into latent variables $z_{t-n}$, which then evolve linearly under the Koopman operator $\mathcal{K}$. The latent variables are subsequently decoded back to approximate the original states. Thus, the latent evolution under Koopman representation captures essential structure but do not directly contain the full state information.
@@
$I(x_{t-1};x_{t})-I(z_{t-1};x_{t})=\underbrace{I(z_{t-1};x_{t}|z_{t})+I(x_{t-1};x_{t}|z_{t-1})}_{\text{step-wise information gap in Koopman representation}}.$ (14)
@@ F.2 Derivation of Variational Distribution
For ergodic systems, conditioning on $x_{0}$ and then taking the long-time average
@@ F.5 Proof of Proposition 1
According to Karush–Kuhn–Tucker (KKT) condition (Boyd & Vandenberghe, 2004), we obtain
@@ G.3.1 Physical Simulation
where $Re=UL/\nu$ is the Reynolds number, defined with characteristic velocity $U$, length $L$, and kinematic viscosity $\nu$. The training dataset covers flows with $Re\in[40,1000]$, while the test dataset focuses on $Re=1000$. The flow is simulated around a cylinder, producing the characteristic alternating vortex shedding pattern. The domain is discretized on a $64\times 64$ grid, with time step $\Delta t=0.001\,\text{s}$, and both $u$ and $v$ velocity components are recorded at each grid point (data from (Takamoto et al., 2022)).
@@
The dam flow dataset is also generated from the two-dimensional incompressible Navier–Stokes equations, using the same formulation as in the Kármán vortex case. The training dataset spans $Re\in[40,1000]$ and the test dataset uses $Re=1000$. The flow is initialized in a rectangular channel with a fixed dam obstacle, where an imposed inlet velocity drives the fluid past the dam-like structure, generating a simple wake pattern downstream. The domain is discretized on a $64\times 64$ spatial grid, with temporal resolution $\Delta t=0.1\,\text{s}$ (data from (Takamoto et al., 2022)). Illustrations are provided in Figure 9.
Figure 9: Examples of physical simulation: Lorenz 63 (left), Kármán Vortex (middle), Dam flow (right).
@@ G.3.2 Visual Inputs
In this task the main goal is to navigate an agent in a surrounded area on a 2D plane (Breivik & Fossen, 2005), whose goal is to navigate from a corner to the opposite one, while avoiding the six obstacles in this area. The system is observed through a set of $40\times 40$ pixel images taken from the top view, which specifies the agent’s location in the area. Actions are two-dimensional and specify the xyx-y direction of the agent’s movement, and given these actions the next positional state of the agent is generated by a deterministic underlying (unobservable) state evolution function. Start State: one of three corners (excluding bottom-right). Goal State: bottom-right corner. Agent’s Objective: agent is within Euclidean distance of 2 from the goal state.
@@
This is the visual version of the classic task of controlling a cart-pole system (Geva & Sitte, 1993). The goal in this task is to balance a pole on a moving cart, while the cart avoids hitting the left and right boundaries. The control (action) is 1-dimensional, which is the force applied to the cart. The underlying state of the system $s_{t}$ is 4-dimensional, which indicates the angle and angular velocity of the pole, as well as the position and velocity of the cart. Similar to the inverted pendulum, in order to maintain the Markovian property the observation $x_{t}$ is a stack of two $80\times 80$ pixel images generated from consecutive time-frames. Start State: Pole is randomly sampled in $\pm\pi/6$. Agent’s Objective: pole’s angle is within $\pm\pi/10$ from an upright position.
@@ G.4 Implementation Algorithm of Three Tasks
Physical Simulation AE (2) $z_{t+1}=\mathcal{K}z_{t}$; reconstruction; Koopman linear forward; InfoNCE; von Neumann entropy
Visual Inputs (Control Tasks) VAE (1) $z_{t+1}=\mathcal{K}z_{t}+Ba_{t}+\epsilon$ (Linear Gaussian); VAE ELBO; reconstruction; InfoNCE; von Neumann entropy
Graph-structured Dynamics AE (2) $z_{t+1}=\mathcal{K}(A)z_{t}$ (adjacency-conditioned); reconstruction; InfoNCE; von Neumann entropy
@@ G.5.4 Model architecture
Table 8: Model architecture for physical simulation with input dimension $(C,H,W)$ with $C$ velocity components and spatial resolution $H\times W$.
@@
Flatten $128C,(\frac{H}{16},\frac{W}{16})\rightarrow\frac{CHW}{4}$
Fully Connected 1 $\frac{CHW}{4}\rightarrow d_{s}$
@@
Decoder Fully Connected 1 $d_{s}\rightarrow\frac{CHW}{4}$ ReLU
Transpose $\frac{CHW}{4}\rightarrow(128C,\frac{H}{16},\frac{W}{16})$
@@
Table 9: Hyperparameter settings for the three tasks in physical simulation
Parameter Symbol Value
temporal coherence $\alpha$ 2.00
von Neumann entropy $\gamma$ 0.10
InfoNCE neighborhood size $k$ 3
Table 10: Hyperparameter settings for the visual inputs
Parameter Symbol Value
temporal coherence $\alpha$ 3.00
@@
von Neumann entropy $\gamma$ 0.5
InfoNCE neighborhood size $k$ 5
Table 11: Hyperparameter settings for graph-structured dynamics
Parameter Symbol Value
temporal coherence $\alpha$ 2.00
von Neumann entropy $\gamma$ 0.10
InfoNCE neighborhood size $k$ 5

### v2

@@ Appendix D Limitations and Future Directions
A current limitation of our framework is that it does not address the sample complexity or non-asymptotic convergence of the Koopman representation; future work could explore more rigorous theoretical analyses in this direction. In addition, recent studies have highlighted connections between kernel methods (Kostic et al., 2022; 2023a; 2024) and information theory (Bach, 2022), suggesting an interesting avenue for extending conventional kernel techniques in Koopman theory through an information-theoretic perspective.
@@ Appendix E Key Technical Definitions and Related Properties
Definition E.2 (Effective Dimension (Roy and Vetterli, 2007))
@@ Appendix F Theoretical Framework
Figure 8: The autoregressive structure of Koopman representation. Top row (solid arrows): the original states $x_{t-n}$ evolve under the nonlinear map $T$. Bottom row (dashed arrows): the states are first encoded into latent variables $z_{t-n}$, which then evolve linearly under the Koopman operator $\mathcal{K}$. The latent variables are subsequently decoded back to approximate the original states. Thus, the latent evolution under Koopman representation captures essential structure but do not directly contain the full state information. Note: the dashed arrows represent the approximated Koopman representation in latent space, whereas the solid arrows denote the true underlying dynamics.
@@
$I(x_{t-1};x_{t})-I(z_{t-1};z_{t})=\underbrace{I(z_{t-1};x_{t}|z_{t})+I(x_{t-1};x_{t}|z_{t-1})}_{\text{step-wise information gap in Koopman representation}}.$ (14)
@@ F.2 Derivation of Variational Distribution
A special case of Proposition 3.1 is ergodic system, conditioning on $x_{0}$ and then taking the long-time average
@@ F.5 Proof of Proposition 1
According to Karush–Kuhn–Tucker (KKT) condition (Boyd and Vandenberghe, 2004), we obtain
@@ G.3.1 Physical Simulation
where $Re=UL/\nu$ is the Reynolds number, defined with characteristic velocity $U$, length $L$, and kinematic viscosity $\nu$. The training dataset covers flows with $Re\in[40,1000]$, while the test dataset focuses on $Re=1000$. The flow is simulated around a cylinder, producing the characteristic alternating vortex shedding pattern. The domain is discretized on a $64\times 64$ grid, with time step $\Delta t=0.001\,\text{s}$, and both $u$ and $v$ velocity components are recorded at each grid point (data from (Yining et al., 2023)).
@@
The dam flow dataset is also generated from the two-dimensional incompressible Navier–Stokes equations, using the same formulation as in the Kármán vortex case. The training dataset spans $Re\in[40,1000]$ and the test dataset uses $Re=1000$. The flow is initialized in a rectangular channel with a fixed dam obstacle, where an imposed inlet velocity drives the fluid past the dam-like structure, generating a simple wake pattern downstream. The domain is discretized on a $64\times 64$ spatial grid, with temporal resolution $\Delta t=0.1\,\text{s}$ (data from (Yining et al., 2023)).
The ERA5 dataset is a global atmospheric reanalysis produced by the European Centre for Medium-Range Weather Forecasts (ECMWF), providing a physically consistent estimate of the large-scale circulation from 1940 to the present (Hersbach et al., 2020). The physic state consists of five channels: 500,hPa geopotential, 850,hPa temperature, 700,hPa specific humidity, and 850,hPa wind components in the zonal and meridional directions. We train all baselines from 1979-01-01 to 2016-01-01 and test after 2018-01-01 (data from (Rasp et al., 2024)). Illustrations are provided in Figure 10.
Figure 10: Examples of physical simulation: (a).Lorenz 63, (b).Kármán Vortex, (c).Dam Flow, (d).ERA5
@@ G.3.2 Visual Inputs
In this task the main goal is to navigate an agent in a surrounded area on a 2D plane (Breivik and Fossen, 2008), whose goal is to navigate from a corner to the opposite one, while avoiding the six obstacles in this area. The system is observed through a set of $40\times 40$ pixel images taken from the top view, which specifies the agent’s location in the area. Actions are two-dimensional and specify the xyx-y direction of the agent’s movement, and given these actions the next positional state of the agent is generated by a deterministic underlying (unobservable) state evolution function. Start State: one of three corners (excluding bottom-right). Goal State: bottom-right corner. Agent’s Objective: agent is within Euclidean distance of 2 from the goal state.
@@
This is the visual version of the classic task of controlling a cart-pole system (Geva and Sitte, 1993). The goal in this task is to balance a pole on a moving cart, while the cart avoids hitting the left and right boundaries. The control (action) is 1-dimensional, which is the force applied to the cart. The underlying state of the system $s_{t}$ is 4-dimensional, which indicates the angle and angular velocity of the pole, as well as the position and velocity of the cart. Similar to the inverted pendulum, in order to maintain the Markovian property the observation $x_{t}$ is a stack of two $80\times 80$ pixel images generated from consecutive time-frames. Start State: Pole is randomly sampled in $\pm\pi/6$. Agent’s Objective: pole’s angle is within $\pm\pi/10$ from an upright position.
@@ G.4 Implementation Algorithm of Three Tasks
Physical Simulation AE (G.2) $z_{t+1}=\mathcal{K}z_{t}$; reconstruction; Koopman linear forward; InfoNCE; von Neumann entropy
Visual Inputs (Control Tasks) VAE (G.1) $z_{t+1}=\mathcal{K}z_{t}+Ba_{t}+\epsilon$ (Linear Gaussian); VAE ELBO; reconstruction; InfoNCE; von Neumann entropy
Graph-structured Dynamics AE (G.2) $z_{t+1}=\mathcal{K}(A)z_{t}$ (adjacency-conditioned); reconstruction; InfoNCE; von Neumann entropy
@@ G.5.1 Physical Simulations
Table 6: Per-channel performance comparison on ERA5 weather forecasting. NN-NRMSE and NN-SSIM denote errors at NN prediction steps; values in parentheses indicate the standard deviation across test samples. Even under the highly stochastic and high-dimensional ERA5 weather dynamics, our approach outperforms all baselines across both short-term and long-term prediction horizons.
Channel Metric KAE KKR PFNN Ours
Geopotential 5-NRMSE 0.058 (0.020) 0.061 (0.027) 0.046 (0.012) 0.023 (0.005)
10-NRMSE 0.068 (0.018) 0.074 (0.026) 0.062 (0.018) 0.032 (0.010)
50-NRMSE 0.157 (0.071) 0.082 (0.025) 0.082 (0.016) 0.075 (0.017)
5-SSIM 0.860 (0.054) 0.852 (0.064) 0.882 (0.036) 0.964 (0.011)
10-SSIM 0.836 (0.051) 0.820 (0.065) 0.848 (0.052) 0.943 (0.028)
50-SSIM 0.665 (0.195) 0.790 (0.049) 0.765 (0.055) 0.806 (0.039)
Temperature 5-NRMSE 0.049 (0.019) 0.052 (0.032) 0.040 (0.009) 0.022 (0.003)
10-NRMSE 0.056 (0.017) 0.064 (0.028) 0.051 (0.015) 0.026 (0.006)
50-NRMSE 0.114 (0.054) 0.074 (0.030) 0.067 (0.018) 0.063 (0.019)
5-SSIM 0.866 (0.046) 0.862 (0.058) 0.888 (0.026) 0.956 (0.008)
10-SSIM 0.844 (0.044) 0.829 (0.063) 0.859 (0.042) 0.942 (0.019)
50-SSIM 0.671 (0.216) 0.802 (0.059) 0.803 (0.051) 0.825 (0.042)
Humidity 5-NRMSE 0.064 (0.029) 0.069 (0.043) 0.055 (0.011) 0.031 (0.003)
10-NRMSE 0.077 (0.027) 0.085 (0.041) 0.070 (0.020) 0.038 (0.006)
50-NRMSE 0.165 (0.086) 0.093 (0.040) 0.088 (0.021) 0.084 (0.028)
5-SSIM 0.859 (0.056) 0.856 (0.076) 0.880 (0.026) 0.954 (0.008)
10-SSIM 0.818 (0.064) 0.805 (0.095) 0.835 (0.053) 0.933 (0.021)
50-SSIM 0.663 (0.220) 0.781 (0.087) 0.776 (0.058) 0.799 (0.057)
Wind $u$ direction 5-NRMSE 0.055 (0.009) 0.056 (0.011) 0.053 (0.006) 0.032 (0.006)
10-NRMSE 0.059 (0.008) 0.061 (0.009) 0.060 (0.008) 0.041 (0.010)
50-NRMSE 0.096 (0.030) 0.063 (0.006) 0.070 (0.007) 0.062 (0.005)
5-SSIM 0.505 (0.122) 0.502 (0.141) 0.537 (0.084) 0.814 (0.051)
10-SSIM 0.433 (0.115) 0.415 (0.134) 0.424 (0.120) 0.721 (0.107)
50-SSIM 0.300 (0.251) 0.361 (0.085) 0.267 (0.100) 0.382 (0.068)
Wind $v$ direction 5-NRMSE 0.051 (0.008) 0.051 (0.009) 0.050 (0.006) 0.031 (0.005)
10-NRMSE 0.054 (0.007) 0.054 (0.008) 0.055 (0.007) 0.040 (0.010)
50-NRMSE 0.060 (0.006) 0.057 (0.005) 0.087 (0.025) 0.057 (0.005)
5-SSIM 0.240 (0.159) 0.247 (0.183) 0.300 (0.091) 0.649 (0.098)
10-SSIM 0.163 (0.133) 0.163 (0.150) 0.208 (0.101) 0.499 (0.163)
50-SSIM 0.105 (0.077) 0.093 (0.082) 0.165 (0.193) 0.094 (0.074)
Table 7: Training time statistics for different models and tasks. Epoch times are reported as mean $\pm$ std (in seconds). For Ours, InfoNCE and entropy (von Neumann entropy) rows correspond to the total per-epoch computation. Notably, the overhead introduced by InfoNCE and von Neumann entropy is marginal, accounting for only a small percentage of the total training time.
Task Metric VAE KAE KKR PFNN Ours
Kármán vortex Epoch time (s) $172.94\pm 4.11$ $195.26\pm 3.47$ $186.47\pm 1.86$ $182.52\pm 2.53$ $201.23\pm 1.07$
InfoNCE time (s) $13.78\pm 0.93$
Entropy time (s) $0.97\pm 0.13$
Dam Flow Epoch time (s) $16.21\pm 0.29$ $16.95\pm 0.33$ $17.48\pm 0.29$ $18.92\pm 0.36$
InfoNCE time (s) $0.76\pm 0.04$
Entropy time (s) $0.56\pm 0.08$
ERA5 Epoch time (s) $40.20\pm 0.52 22$$24.95\pm 0.54$ $242.33\pm 0.71$ $253.24\pm 2.05$
InfoNCE time (s) $15.09\pm 0.80$
Entropy time (s) $7.43\pm 0.64$
Figure 16: Comparison of continuous predictions for the global geopotential starting from 2018010100:002018-01-01-00:$0$ to 2018010808:002018-01-08-08:$0$. Ground truth are contrasted with predictions from KAE, KKR, PFNN, and our method. Error maps in the lower panels demonstrate that, compared with other models, showing with more stable and accurate results of our model.
Figure 17: Comparison of continuous predictions for the global temperature starting from 2018010100:002018-01-01-00:$0$ to 2018010808:002018-01-08-08:$0$. Ground truth are contrasted with predictions from KAE, KKR, PFNN, and our method. Error maps in the lower panels demonstrate that, compared with other models, showing with more stable and accurate results of our model.
Figure 18: Comparison of continuous predictions for the global $u$-direction wind starting from 2018010100:002018-01-01-00:$0$ to 2018010808:002018-01-08-08:$0$. Ground truth are contrasted with predictions from KAE, KKR, PFNN, and our method. Error maps in the lower panels demonstrate that, compared with other models, showing with more stable and accurate results of our model.
Figure 19: Comparison of continuous predictions for the global $v$-direction wind starting from 2018010100:002018-01-01-00:$0$ to 2018010808:002018-01-08-08:$0$. Ground truth are contrasted with predictions from KAE, KKR, PFNN, and our method. Error maps in the lower panels demonstrate that, compared with other models, showing with more stable and accurate results of our model.
Figure 20: Comparison between $\gamma=0.0$ and $\gamma=0.5$ on the physical simulation task with latent dimension $32$. (a) Evolution of the eigenvalue spectrum over training epochs. (b) Heatmap of the latent covariance matrix for $\gamma=0.0$ across epochs. (c) Heatmap of the latent covariance matrix for $\gamma=0.5$ across epochs. With the addition of the von Neumann entropy regularizer, two clear effects emerge during training. First, the latent covariance matrix no longer collapses to a few dominant modes: the eigenvalue distribution becomes more uniform and remains close to full rank throughout optimization (see (a)), indicating that the model learns a richer set of modes rather than compressing them into a low-dimensional subspace. Second, the covariance structure transitions from highly sparse (when $\gamma=0$) to dense and full-rank under entropy regularization (from (b) to (c)), verifying our theoretical prediction in Proposition 1.
Figure 21: Visualization of the von Neumann entropy regularization loss and the total training loss over epochs for the physical simulation tasks. The stable behavior of both the total loss and the von Neumann entropy loss indicates that our training procedure is numerically stable and robust across different systems.
@@ G.5.4 Model architecture
Table 10: Model architecture for Kármán Vortex and Dam Flow task with input dimension $(C,H,W)$, where $C$ denotes the number of velocity components and $H\times W$ is the spatial resolution.
@@
Flatten $128C,(\frac{H}{16},\frac{W}{16})\rightarrow\frac{CHW}{2}$
Fully Connected 1 $\frac{CHW}{2}\rightarrow d_{s}$
@@
Decoder Fully Connected 1 $d_{s}\rightarrow\frac{CHW}{2}$ ReLU
Transpose $\frac{CHW}{2}\rightarrow(128C,\frac{H}{16},\frac{W}{16})$
@@
Table 11: Model architecture for ERA5 task with input dimension $(C,H,W)$, where $C$ denotes the number of channels and $H\times W$ is the spatial resolution, using a factorized-attention encoder.
Components Layer Layer number $C,(H,W)$ Activation
Encoder Conv2d 1 $C\rightarrow\frac{64C}{5},\ (H,W)$
Conv2d 1 $\frac{64C}{5}\rightarrow\frac{64C}{5},\ (\frac{H}{4},\frac{W}{4})$
FactorizedBlock 1 $\frac{64C}{5},\ (\frac{H}{4},\frac{W}{4})$ GELU
Flatten $\frac{64C}{5},\ (\frac{H}{4},\frac{W}{4})\rightarrow\frac{4CHW}{5}$
Fully Connected 1 $\frac{4CHW}{5}\rightarrow d_{s}$
Koopman Operator Linear 1 $d_{s}\rightarrow d_{s}$
Decoder Fully Connected 1 $d_{s}\rightarrow\frac{4CHW}{5}$ ReLU
Transpose $\frac{4CHW}{5}\rightarrow(\frac{64C}{5},\frac{H}{4},\frac{W}{4})$
ConvTranspose2d 2 $\frac{64C}{5}\rightarrow\frac{64C}{5},(H,W)$ ReLU
Conv2d 1 $\frac{64C}{5}\rightarrow C,\ (H,W)$
Conv2d Refinement 3 $C\rightarrow C,\ (H,W)$ ReLU
@@
Table 12: Hyperparameter settings across physical simulations, visual inputs, and graph-structured dynamics
Parameter Symbol Physical Sim. Visual Inputs Graph Dyn.
Temporal coherence $\alpha$ 2.00 3.00 2.00
@@
von Neumann entropy $\gamma$ 0.10 0.50 0.10
InfoNCE neighborhood $k$ 3 5 5